# Identification of gapless phases by squaring a twist operator

### v1

@@ Appendix A Supplemental Materials
A.1 Analytical results of $\langle\Psi_{\text{M-G}}|\hat{U}^{2}|\Psi_{\text{M-G}}\rangle$ for Majumdar-Ghosh model defined on even-length chains
As stated in the text, there are two degenerated VBS ground states when the length is chosen as an even number $L\in 2\mathbb{Z}$,
@@
$\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle$ $=$ $\left[\cos\left(\frac{2\pi}{L}\right)\right]^{L/2},$ (30)
$\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle$ $=$ $\left[\cos\left(\frac{2\pi}{L}\right)\right]^{L/2-1}\cos\left[\frac{2\pi}{L}(L-1)\right].$ (31)
By taking thermodynamic limit $L\rightarrow\infty$, we obtain
$\lim_{L\rightarrow\infty}\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle$ $=$ $1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}-\frac{4\pi^{4}+\pi^{6}}{6L^{3}}+\mathscr{O}\left(\frac{1}{L^{4}}\right),$ (32)
$\lim_{L\rightarrow\infty}\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle$ $=$ $1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}-\frac{(-2/3)\pi^{4}+(1/6)\pi^{6}}{6L^{3}}+\mathscr{O}\left(\frac{1}{L^{4}}\right).$ (33)
@@
$\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle\propto\left(\frac{1}{\sqrt{2}}\right)^{L/2}=\mathscr{O}\left(\frac{1}{L^{+\infty}}\right).$ (34)
We observe that the coefficients of $1/L$ and $1/L^{2}$ terms are the same for these two VBS ground states, resulting in the same coefficients of leading term for an arbitrary normalized ground state $|\Psi_{\text{M-G}}\rangle\equiv\alpha|\Psi_{\text{M-G}}^{(1)}\rangle+\beta|\Psi_{\text{M-G}}^{(2)}\rangle$ spanned by them, with $|\alpha|^{2}+|\beta|^{2}=1$,
@@
$\lim_{L\rightarrow\infty}\langle\Psi_{\text{M-G}}|\hat{U}^{2}|\Psi_{\text{M-G}}\rangle$ $=$ $\lim_{L\rightarrow\infty}\left(|\alpha|^{2}\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle+|\beta|^{2}\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle\right)$ (35)
$=$ $\left(|\alpha|^{2}+|\beta|^{2}\right)\left[1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}+\mathscr{O}\left(\frac{1}{L^{3}}\right)\right]$
$=$ $1-\frac{\pi^{2}}{L}+\frac{\pi^{4}}{2L^{2}}+\mathscr{O}\left(\frac{1}{L^{3}}\right).$

### v2

@@ Appendix A Supplemental Materials
Appendix B Analytical results of $\langle\hat{U}^{2}\rangle$ for Majumdar-Ghosh model defined on even-length chains
B.1 Periodic boundary conditions (PBCs) results
As stated in the text, there are two degenerated valence-bond-state (VBS) type ground states when the length is chosen as an even number $L\in 2\mathbb{Z}$ under PBCs,
@@ B.1 Periodic boundary conditions (PBCs) results
$\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle$ $=$ $\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle=\left[\cos\left(\frac{2\pi}{L}\right)\right]^{L/2}=1-\frac{\pi^{2}}{L}+\mathscr{O}(1/L^{2}).$ (30)
@@
$\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle\propto\left(\frac{1}{\sqrt{2}}\right)^{L/2}=\mathscr{O}\left(\frac{1}{L^{+\infty}}\right);$
@@
$\langle\Psi_{\text{M-G}}^{(2)}|\Psi_{\text{M-G}}^{(1)}\rangle\propto\left(\frac{1}{\sqrt{2}}\right)^{L/2}=\mathscr{O}\left(\frac{1}{L^{+\infty}}\right).$ (31)
Therefore, for an arbitrary normalized ground state $|\Psi_{\text{M-G}}\rangle\equiv\alpha|\Psi_{\text{M-G}}^{(1)}\rangle+\beta|\Psi_{\text{M-G}}^{(2)}\rangle$ spanned by them, with $|\alpha|^{2}+|\beta|^{2}=1+\mathscr{O}(1/L^{\infty})$,
$\langle\Psi_{\text{M-G}}|\hat{U}^{2}|\Psi_{\text{M-G}}\rangle$ $=$ $\left(|\alpha|^{2}\langle\Psi_{\text{M-G}}^{(1)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(1)}\rangle+|\beta|^{2}\langle\Psi_{\text{M-G}}^{(2)}|\hat{U}^{2}|\Psi_{\text{M-G}}^{(2)}\rangle\right)$ (32)
$=$ $\left(|\alpha|^{2}+|\beta|^{2}\right)\left[1-\frac{\pi^{2}}{L}+\mathscr{O}\left(\frac{1}{L^{2}}\right)\right]$
$=$ $1-\frac{\pi^{2}}{L}+\mathscr{O}\left(\frac{1}{L^{2}}\right).$
B.2 Open Boundary Conditions (OBCs) results
In the presence of OBCs, the lowest energy states becomes 5-fold degenerate:
$|\psi_{\text{M-G}}\rangle$ $=$ $|(1,2)(3,4)\ldots(L-1,L)\rangle,$ (33)
$|\psi_{\text{M-G}}^{(i,j)}\rangle$ $=$ $|(-1)^{i+1},(2,3)(4,5)\ldots(L-2,L-1),(-1)^{j+1}\rangle,\,\,(i,j)\in\{1,2\},$ (34)
where there are 4 states with dangling edge spin-$1/2$ excitations. Calculation of $\langle\hat{U}^{2}\rangle$ of these five degenerate states satisfies the same rule as PBCs:
$\langle\psi_{\text{M-G}}|\hat{U}^{2}|\psi_{\text{M-G}}\rangle$ $=$ $1-\frac{\pi^{2}}{L}+\mathscr{O}\left(\frac{1}{L^{2}}\right);$
$\langle\psi_{\text{M-G}}^{(i,j)}|\hat{U}^{2}|\psi_{\text{M-G}}^{(i,j)}\rangle$ $=$ $\left(\begin{array}[]{cc}1-\frac{\pi^{2}}{L}+i\frac{2\pi}{L}&1-\frac{\pi^{2}}{L}+i\frac{2\pi}{L}\\ 1-\frac{\pi^{2}}{L}-i\frac{2\pi}{L}&1-\frac{\pi^{2}}{L}-i\frac{2\pi}{L}\end{array}\right)_{i,j}+\mathcal{O}(1/L^{2})$ (37)
$=$ $1+\mathcal{O}(1/L).$ (38)
Nevertheless, for a general linear combination, e.g., two special cases as
$\left.\begin{array}[]{c}\left[\frac{1}{\sqrt{2}}\left(\langle\psi_{\text{M-G}}^{(1,2)}|\pm\langle\psi_{\text{M-G}}^{(2,1)}|\right)\right]\hat{U}^{2}\left[\frac{1}{\sqrt{2}}\left(|\psi_{\text{M-G}}^{(1,2)}\rangle\pm|\psi_{\text{M-G}}^{(2,1)}\rangle\right)\right]\\ \left[\frac{1}{\sqrt{2}}\left(\langle\psi_{\text{M-G}}^{(1,1)}|\pm\langle\psi_{\text{M-G}}^{(2,2)}|\right)\right]\hat{U}^{2}\left[\frac{1}{\sqrt{2}}\left(|\psi_{\text{M-G}}^{(1,1)}\rangle\pm|\psi_{\text{M-G}}^{(2,2)}\rangle\right)\right]\end{array}\right\}=1-\frac{\pi^{2}}{L}+\mathcal{O}(1/L^{2}),$ (41)
which means that the coefficients of $1/L$ are not the same for a general linear combinations, unlike PBCs cases before. It is not a coincidence that there is no imaginary part in Eq. (93) as well as all the PBCs case; all these states are eigenstates of the global $\pi$-rotation transformation around $x$ axis, so $\langle\hat{U}^{2}\rangle$ must be real.
Appendix C Analytical results of $\langle\hat{U}^{2}\rangle$ for Affleck-Kennedy-Lieb-Tasaki model
C.1 PBCs results
We consider the ground state of AKLT model in translation invariant matrix product states (MPS) form under PBCs [48, 36, 49]:
$\langle\bm{\sigma}|\Psi_{\text{AKLT}}\rangle$ $=$ $\sum_{\alpha_{1},\alpha_{2},\cdots,\alpha_{L}\in\{1,2\}}A_{\alpha_{1},\alpha_{2}}^{\sigma_{1}}A_{\alpha_{2},\alpha_{3}}^{\sigma_{2}}\cdots A_{\alpha_{L},\alpha_{1}}^{\sigma_{L}}$ (42)
$=$ $\mathrm{Tr}[\bm{A}^{\sigma_{1}}\bm{A}^{\sigma_{2}}\cdots\bm{A}^{\sigma_{L}}],$ (43)
where the vector $\bm{\sigma}=\{\sigma_{1},\sigma_{2},\cdots,\sigma_{L}\}$ represents the physical spin-$1$ configuration, and it can be defined for matrices $\bm{A}^{\sigma}$ in the second line in the $\{0,\pm\}$-basis as
$\bm{A}^{+}=\left(\begin{array}[]{cc}0&0\\ -\frac{1}{\sqrt{2}}&0\end{array}\right),\quad\bm{A}^{0}=\left(\begin{array}[]{cc}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{array}\right),\quad\bm{A}^{-}=\left(\begin{array}[]{cc}0&\frac{1}{\sqrt{2}}\\ 0&0\end{array}\right).$ (50)
We note that the VBS state is not normalized, which can be calculated out
$\langle\Psi_{\text{AKLT}}|\Psi_{\text{AKLT}}\rangle=\mathrm{Tr}[\tilde{\bm{A}}^{L}]=\left(\frac{3}{4}\right)^{L}+3\left(-\frac{1}{4}\right)^{L},$ (51)
where
$\tilde{\bm{A}}=\left(\begin{array}[]{cccc}\frac{1}{4}&\frac{1}{2}&0&0\\ \frac{1}{2}&\frac{1}{4}&0&0\\ 0&0&-\frac{1}{4}&0\\ 0&0&0&-\frac{1}{4}\end{array}\right).$ (56)
For the twisting operator, we consider the general case for $\langle\hat{U}^{q}\rangle$ for any integer $q$,
$\hat{U}^{q}=\exp\left(\frac{2\pi qi}{L}\sum_{j=1}^{L}j\hat{S}^{z}_{j}\right),$ (57)
and define a site-dependent parameter for the site-$j$
$\theta^{(j)}=2\pi qj/L.$ (58)
Then, one can yield out for spin-1 system,
$e^{i\theta^{(j)}\hat{S}_{j}^{z}}=\mathrm{diag}\{e^{i\theta^{(j)}},1,e^{-i\theta^{(j)}}\}.$ (59)
Finally, we deduce out the expectation value for the general twisting operator $\hat{U}^{q}$
$\langle\hat{U}^{q}\rangle=\mathrm{Tr}\left[\prod_{j=1}^{L}\tilde{\bm{C}}^{(j)}\right]\bigg{/}\langle\Psi_{\text{AKLT}}|\Psi_{\text{AKLT}}\rangle,$ (60)
where
$\tilde{\bm{C}}^{(j)}=\left(\begin{array}[]{cccc}\frac{1}{4}&\frac{1}{2}e^{i\theta^{(j)}}&0&0\\ \frac{1}{2}e^{-i\theta^{(j)}}&\frac{1}{4}&0&0\\ 0&0&\frac{1}{4}&0\\ 0&0&0&\frac{1}{4}\end{array}\right).$ (65)
We can further simplify it to
$\langle\hat{U}^{q}\rangle$ $=$ $\left\{\mathrm{Tr}\left[\prod_{j=1}^{L}\tilde{\bm{D}}^{(j)}\right]+\mathrm{Tr}\left[\prod_{j=1}^{L}\tilde{\bm{E}}^{(j)}\right]\right\}\bigg{/}\langle\Psi_{\text{AKLT}}|\Psi_{\text{AKLT}}\rangle$ (66)
$=$ $\left\{\mathrm{Tr}\left[\prod_{j=1}^{L}\tilde{\bm{D}}^{(j)}\right]+\left(\frac{2}{4}\right)^{L}\right\}\bigg{/}\left[\left(\frac{3}{4}\right)^{L}+3\left(-\frac{1}{4}\right)^{L}\right]$
$=$ $\left\{\mathrm{Tr}\left[\prod_{j=1}^{L}{\bm{D}}^{(j)}\right]+\frac{2^{L}}{3^{L}}\right\}\bigg{/}\left[1-\left(-\frac{1}{3}\right)^{L-1}\right],$
where
$\tilde{\bm{D}}^{(j)}=\left(\begin{array}[]{cccc}\frac{1}{4}&\frac{1}{2}e^{i\theta^{(j)}}\\ \frac{1}{2}e^{-i\theta^{(j)}}&\frac{1}{4}\end{array}\right),\ \ \ \tilde{\bm{E}}^{(j)}=\left(\begin{array}[]{cccc}\frac{1}{4}&0\\ 0&\frac{1}{4}\end{array}\right),$ (71)
and
$\bm{D}^{(j)}=\frac{1}{3}\left(\begin{array}[]{cccc}1&2e^{i\theta^{(j)}}\\ 2e^{-i\theta^{(j)}}&1\end{array}\right).$ (74)
When $q=2$, the trace of product of $\bm{D}^{(j)}$ is the leading term, then
$\mathrm{Tr}\left[\prod_{j=1}^{L}\bm{D}^{(j)}\right]$ $=$ $\mathrm{Tr}\left[\frac{1}{3}\left(\begin{array}[]{cccc}1&2\\ 2&1\end{array}\right)\exp\left(i\sigma_{z}\frac{2\pi}{L}\right)\right]^{L}$
$=$ $\frac{1}{3^{L}}\sum_{\eta=\pm}\left(\cos\frac{2\pi}{L}+\eta\sqrt{4-\sin^{2}\frac{2\pi}{L}}\right)^{L}$
$=$ $1-\frac{\pi^{2}}{L}+\mathscr{O}(1/L^{2}).$ (78)
C.2 OBCs results
Under OBCs, AKLT model has 4-fold degenerate ground states [48, 36, 49]:
$\langle\bm{\sigma}|\Psi^{(i,j)}_{\text{AKLT}}\rangle$ $=$ $\sum_{\alpha_{1},\alpha_{2},\cdots,\alpha_{L}\in\{1,2\}}A_{i,\alpha_{2}}^{\sigma_{1}}A_{\alpha_{2},\alpha_{3}}^{\sigma_{2}}\cdots A_{\alpha_{L},j}^{\sigma_{L}}$ (79)
$=$ $(\bm{A}^{\sigma_{1}}\bm{A}^{\sigma_{2}}\cdots\bm{A}^{\sigma_{L}})_{i,j},$ (80)
where $i,j\in\{1,2\}$. These state has spin-$1/2$ magnetization $(-1)^{i+1}/2$ and $(-1)^{j+1}/2$ around sites $1$ and $L$, respectively.
Their norms can be calculated as
$\langle\Psi^{(i,j)}_{\text{AKLT}}|\Psi^{(i,j)}_{\text{AKLT}}\rangle=[\tilde{\bm{A}}^{L}]_{i,j}=\frac{1}{2\times 4^{L}}\left(\begin{array}[]{cc}3^{L}+(-1)^{L}&3^{L}-(-1)^{L}\\ 3^{L}-(-1)^{L}&3^{L}+(-1)^{L}\end{array}\right)_{i,j},$ (83)
and then, for a fixed ground state labelled by $(i,j)\in\{1,2\}$:
$\langle\Psi^{(i,j)}_{\text{AKLT}}|\hat{U}^{2}|\Psi^{(i,j)}_{\text{AKLT}}\rangle$ $=$ $\left[\prod_{k=1}^{L}\tilde{\bm{C}}^{(k)}\right]_{i,j}\bigg{/}\langle\Psi^{(i,j)}_{\text{AKLT}}|\Psi^{(i,j)}_{\text{AKLT}}\rangle$ (86)
$=$ $\frac{1}{4^{L}}\left[\left(\begin{array}[]{cccc}1&2\\ 2&1\end{array}\right)\exp\left(i\sigma_{z}\frac{2\pi}{L}\right)\right]^{L}{}_{i,j}\bigg{/}\langle\Psi^{(i,j)}_{\text{AKLT}}|\Psi^{(i,j)}_{\text{AKLT}}\rangle$
$=$ $\left(\begin{array}[]{cc}1-\frac{\pi^{2}}{L}+i\frac{\pi}{L}&1-\frac{\pi^{2}}{L}-i\frac{2\pi}{L}\\ 1-\frac{\pi^{2}}{L}+i\frac{2\pi}{L}&1-\frac{\pi^{2}}{L}-i\frac{\pi}{L}\end{array}\right)_{i,j}+\mathcal{O}(1/L^{2})$ (89)
$=$ $1+\mathcal{O}(1/L),$ (90)
where we can see that the asymptotical rule $\mathcal{O}(1/L)$ is still preserved under OBCs although the coefficients are modified and not the same for a general linear combination.
Similarly as Eq. (93) and consistently with the global $\pi$-rotation around $x$-axis:
$\left.\begin{array}[]{c}\left[\frac{1}{\sqrt{2}}\left(\langle\Psi_{\text{AKLT}}^{(1,2)}|\pm\langle\Psi_{\text{AKLT}}^{(2,1)}|\right)\right]\hat{U}^{2}\left[\frac{1}{\sqrt{2}}\left(|\Psi_{\text{AKLT}}^{(1,2)}\rangle\pm|\Psi_{\text{AKLT}}^{(2,1)}\rangle\right)\right]\\ \left[\frac{1}{\sqrt{2}}\left(\langle\Psi_{\text{AKLT}}^{(1,1)}|\pm\langle\Psi_{\text{AKLT}}^{(2,2)}|\right)\right]\hat{U}^{2}\left[\frac{1}{\sqrt{2}}\left(|\Psi_{\text{AKLT}}^{(1,1)}\rangle\pm|\Psi_{\text{AKLT}}^{(2,2)}\rangle\right)\right]\end{array}\right\}$ $=$ $1-\frac{\pi^{2}}{L}+\mathcal{O}(1/L^{2})\in\mathbb{R}.$ (93)